OpenStudy (anonymous):

TWENTY IS WHAT PERCENT OF 200 A 20% B 10% C40% D80%

OpenStudy (abbles):

10 percent

OpenStudy (anonymous):

what is 20/200?

OpenStudy (anonymous):

yeah, what Abbles said

OpenStudy (abbles):

you would move the decimal point one spot to the left

OpenStudy (anonymous):

i thought is 40

OpenStudy (anonymous):

it is the ratio of 20 to 200, or 20/200 which is the same as 1/10 or 10%

OpenStudy (anonymous):

It is 10%. You can use a proportion to solve it. 20/200 = x/100. Cross multiply, 20*100=200*x. 2000=200x divide by 200 10=x So it is 10%.

OpenStudy (anonymous):

o ya it is 10%